# Evaluating Multimodal Large Language Models on Educational Textbook Question Answering

### v1

@@ Abstract
Multimodal large language models (MLLMs) have recently achieved significant success in vision–language tasks. However, their capacity to reason over complex, long lessons and intricate educational diagrams that cannot be represented as a single natural image, remains largely untested. In this work, we present the first evaluation of state-of-the-art MLLMs on the textbook question answering (TQA) task using the CK12-QA dataset. We assess the performance of recent vision-language models, including LLaVA and LLaMA 3.2-Vision, across various input configurations. Additionally, we introduce a lightweight multimodal retrieval-augmented generation (RAG) pipeline that integrates both paragraphs and diagrams from the lesson into the prompt. Our results demonstrate the influence of retrieved educational context on model accuracy and reasoning, while also revealing current limitations in handling question-context relationships and the potential for noise, pointing to key directions for future research in multimodal AI-driven learning.
@@
Educational AI, Textbook Question Answering, Multimodal Large Language Models, Multimodal Retrieval-Augmented Generation
@@ I Introduction
Answering curriculum-related questions in multimodal educational materials is a central challenge in AI for education, requiring systems to reason across complex multimodal contexts such as lengthy lessons, diagrams, and videos. While general-purpose visual question answering (VQA) has made substantial progress with multimodal large language models (MLLMs) such as LLaVA [1, 2] and LLaMA 3.2-Vision [3], most evaluations have focused on natural images or open-domain reasoning, overlooking the distinct demands of textbook question answering (TQA). TQA requires not only parsing long-form lessons but also interpreting intricate diagrams, integrating supplementary resources such as videos, and analyzing comprehensive—often fragmented—curricular content that exceeds the scope of standard VQA tasks [4]. This underscores the significance of our contribution in addressing the challenges of TQA, particularly in zero-shot academic settings.
@@
The CK12-QA questions illustrated in Fig. 1 are multimodal, multiple-choice science questions designed to engage students in mid-level cognitive tasks, including interpretation, conceptual reasoning, and visual comprehension. Unlike simple fact-recall questions, these items require a solid understanding of scientific concepts, such as the layers of the Earth, airflow over mountains, stages of the cell cycle, and semiconductor behavior, and the ability to synthesize textual prompts with informative diagrams. For example, students must deduce correct answers by analyzing labeled images or interpreting scientific processes. This aligns the questions with mid-tier levels of Bloom’s taxonomy, including understanding, applying, and analyzing; skills that are essential for deeper comprehension in K–12 science education.
@@
Contributions. This paper makes the following contributions:
  • We benchmark two state-of-the-art multimodal large language models (MLLMs), LLava 1.5 and LLaMA 3.2-Vision, on the CK12-QA dataset, evaluating their performance on both diagram-based and text-only questions.
  • We compare model performance across different input configurations, including standalone questions and retrieval-augmented settings, to assess the impact of contextual information.
  • We conduct a quantitative analysis of the models’ ability to align multimodal questions with multimodal context.
  • We propose a retrieval-augmented evaluation framework for textbook question answering (TQA), enabling systematic investigation of context-aware prompting strategies for MLLM-based answer generation.
This work advances our understanding of how MLLMs can be leveraged for complex educational reasoning tasks, highlighting both their potential and their current limitations in supporting multimodal, curriculum-based learning.
@@ III Method
Given a query embedding $f(q)\in\mathbb{R}^{d}$, derived from the question text or image depending on the modality, and a set of candidate embeddings $g(c_{i})\in\mathbb{R}^{d}$ representing paragraphs or diagrams, we compute the cosine similarity to rank candidates, as shown in equation 1.
@@
All embeddings are generated using ImageBind [18], a cross-modal embedding model that projects inputs from various modalities, such as text, images, and audio, into a shared embedding space. To store and index paragraph embeddings, we utilize KDB.AI [19], a high-performance, scalable vector database optimized for similarity search. During inference, the question embedding is used to query KDB.AI, which efficiently retrieves the top-$k$ context based on their cosine similarity to the query.
@@
Each input is composed of the following elements:
  • Question: A multiple-choice science question.
  • (Optional) Associated Diagram: An educational illustration related to the question (e.g., a labeled process or structure).
  • Multimodal Context: A relevant paragraph and diagrams retrieved from the full set of CK12 lessons using a RAG approach.
@@
No architectural modifications are made to underlying models. Instead, the methodology focuses on leveraging existing pretrained MLLMs within a structured TQA framework.
IV Experiments
@@
0p-01, 1p-11, 3p-3I = number of retrieved paragraphs and images; T-T/F = Textual True/False; T-MC = Textual Multiple Choice; All = overall accuracy across DQ and NDQ.
In this section, we present our experimental design, results, and analysis used to evaluate the capabilities of MLLMs in answering curriculum-aligned questions. The experiments are designed to assess performance across varying input modalities and to examine the impact of retrieval-augmented context on structured academic tasks.
IV-A Experimental Setup
@@
IV-A1 Dataset
We utilize the CK12-QA dataset, a curriculum-aligned benchmark comprising both diagram-based and text-based science questions. The dataset includes textual true/false (T-T/F), textual multiple choice (T-MC), and diagram-based (Diagram Qs) items drawn from 1,076 lessons across Life Science, Earth Science, and Physical Science.
IV-A2 Input Conditions
Each model is evaluated under a zero-shot setting using two input configurations:
  • Basic (Image + Question): Only the question and its associated diagram (if applicable) are provided.
  • Multimodal RAG (Image + Question + Context): A retrieved lesson paragraph is included alongside the question and its diagram.
IV-A3 Evaluation Metrics
@@
where $\hat{y}^{(n)}$ is the model prediction, and $\mathbb{1}[\cdot]$ is the indicator.
IV-B Results
Tables III summarize validation and test accuracy for both models across varying input configurations. Key findings include:
Task-Specific Performance
  • Text-Only Questions (T-T/F, T-MC): LLaMA demonstrates stronger baseline performance without additional context (e.g., 74.50% vs. 39.04% on T-MC in the 0p-0I setting–test set). However, LLaVA shows greater relative improvement when context is introduced. In the 3p-3I setting–test set, LLaVA’s T-MC accuracy improves by +14.16% (from 39.04% to 53.20%), narrowing the gap with LLaMA, whose performance declines by –17.11% (from 74.50% to 57.39%) on the test set. This indicates that LLaVA more effectively leverages retrieved textual context, whereas LLaMA’s performance is more sensitive to the presence of distractor paragraphs.
  • Diagram Questions (DQ): LLaMA significantly outperforms LLaVA on baseline diagram-based questions, achieving 74.07% accuracy on the validation set in the 0p-0I setting compared to LLaVA’s 46.89%.
    However, with added context, LLaMA’s performance on diagram questions–validation set degrades sharply, dropping to 25.93% in the 3p-3I setting, while LLaVA remains more stable at 42.89%. This suggests that LLaVA is more robust to multimodal interference introduced by retrieval-augmented input.
@@ IV-B Results
IV-C Fine-Tuned Model Results
During fine-tuning, we minimize the autoregressive cross-entropy loss:
@@ IV-C Fine-Tuned Model Results
TABLE III: Performance of Fine-Tuned LLaMA 3.2-Vision on CK12-QA Dataset under the 3 Paragraph, 3 Image Configuration.
Metric Validation (3p-3I) Test (3p-3I)
@@
TABLE IV: Performance of Fine-Tuned LLaVA on CK12-QA Dataset under the 3 Paragraph, 3 Image Configuration.
Metric Validation (3p-3I) Test (3p-3I)
@@
LLaVA shows a marked decline in overall performance following fine-tuning, with accuracy dropping from 48.92% to 27.78% on the validation set. Notably, the model exhibits inconsistent behavior across tasks in both test and validation sets, for example: while performance on T-T/F questions had a marginal drop on the test set (from 53.71% to 50.98%), accuracy on T-MC questions falls sharply to 19.24% from 53.20%, indicating instability in task generalization. Accuracy on diagram-based questions decreased across both validation and test sets, suggesting limited improvement in visual–textual alignment despite fine-tuning.
Fine-tuning outcomes diverge sharply between the two models. As shown in Table III, LLaMA 3.2-Vision demonstrates substantial improvements over its zero-shot performance, achieving 71.16% accuracy on the test set, up from 35.31%. Significant gains are observed across all task types: for example, on the validation set, T-MC accuracy increases from 59.15% to 79.76%, while performance on diagram-based questions rises from 25.93% to 65.88%, indicating effective multimodal learning. These results reflect strong cross-modal integration and suggest that fine-tuning successfully mitigates context interference present in zero-shot settings. In contrast, LLaVA struggles to adapt effectively, as shown in Table IV. Its validation accuracy drops by nearly 50%, while LLaMA 3.2-Vision achieves 73.72% overall accuracy on the same split. This disparity highlights potential architectural advantages in LLaMA for multimodal fusion. LLaVA also shows task-specific instability: although T-T/F performance a slight decrease on the test and validation sets, T-MC accuracy declines markedly, pointing to inconsistent task prioritization. By comparison, LLaMA maintains balanced improvements across all tasks, suggesting more robust and curriculum-aware reasoning capabilities. Finally, whereas LLaMA effectively leverages retrieved context and resolves conflicts between textual and diagrammatic inputs, LLaVA appears to either overfit to noisy context or lack sufficient visual grounding—evidenced by its persistently low performance on diagram-based questions.
@@
  • LLaVA in the zero-shot setting, shows a consistent ability to leverage retrieved context, with overall validation accuracy increasing from 44.45% in the 0p-0I setting to 48.92% in the 3p-3I setting. This improvement highlights its capacity to integrate additional textual information for enhanced reasoning.
  • LLaMA 3.2-Vision in the zero-shot setting, achieves strong baseline performance on diagram-based questions, with validation accuracy reaching 74.07% in the 0p-0I setting. However, performance drops sharply to 25.93% in the 3p-3I setting when contextual paragraphs are added, indicating difficulty in fusing retrieved text with visual information.
IV-D Discussion
@@
This discrepancy underscores the importance of context relevance filtering. Despite their strong vision-language capabilities, both models exhibit difficultly handling indiscriminately retrieved paragraphs. To mitigate this, more effective context selection strategies are needed—strategies that account not only for textual similarity but also for visual–textual alignment. Incorporating cross-modal attention mechanisms may offer a promising solution, enabling models to dynamically evaluate and prioritize context passages based on their relevance to both the question and associated visual input.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have shown success in vision-language tasks, but their ability to reason over complex educational materials remains largely untested. This work presents the first evaluation of state-of-the-art MLLMs, including LLaVA-1.5 and LLaMA 3.2-Vision, on the textbook question answering (TQA) task using the CK12-QA dataset. We introduce a multimodal retrieval-augmented generation (RAG) pipeline to simulate real-world learning by providing relevant lesson paragraphs and diagrams as context. Our zero-shot experiments reveal a critical trade-off; while retrieved context improves LLaVA’s performance on text-based questions, it significantly degrades the accuracy of the more powerful LLaMA 3.2-Vision on diagram-based tasks, dropping its validation accuracy from 74.07% to 25.93%. We term this statistically significant phenomenon ”catastrophic context interference.” Furthermore, fine-tuning highlights architectural differences; LLaMA 3.2-Vision’s performance substantially improves to 71.16% on the test set, demonstrating its capacity to learn multimodal integration, whereas LLaVA’s performance declines, indicating challenges with generalization. Our results underscore the challenges MLLMs face in modality prioritization and context integration, providing a benchmark and pointing to key directions for developing more robust AI-driven educational tools.
@@
Educational AI, Multimodal Large Language Models, Multimodal Retrieval-Augmented Generation, Textbook Question Answering.
@@ I Introduction
Answering curriculum-related questions in multimodal educational materials is a central challenge in AI for education, requiring systems to reason across complex multimodal contexts such as lengthy lessons, diagrams, and videos. While general-purpose visual question answering (VQA) has made substantial progress with multimodal large language models (MLLMs) such as LLaVA [1, 2] and LLaMA 3.2-Vision [3], most evaluations have focused on natural images or open-domain reasoning, overlooking the distinct demands of textbook question answering (TQA). TQA requires not only parsing long-form lessons but also interpreting intricate diagrams, integrating supplementary resources such as videos, and analyzing comprehensive, often fragmented, curricular content that exceeds the scope of standard VQA tasks [4]. This underscores the significance of our contribution in addressing the challenges of TQA.
@@
The CK12-QA questions illustrated in Fig. 1 are multimodal, multiple-choice science questions designed
@@
to engage students in mid-level cognitive tasks, including interpretation, conceptual reasoning, and visual comprehension. Unlike simple fact-recall questions, these items require a solid understanding of scientific concepts, such as the layers of the Earth, airflow over mountains, stages of the cell cycle, and semiconductor behavior, and the ability to synthesize textual prompts with informative diagrams. For example, students must deduce correct answers by analyzing labeled images or interpreting scientific processes. This aligns the questions with mid-tier levels of Bloom’s taxonomy, including understanding, applying, and analyzing; skills that are essential for deeper comprehension in K–12 science education.
@@
Contributions. This paper makes several key contributions to the field. We first benchmark two state-of-the-art multimodal large language models, LLaVA 1.5 and LLaMA 3.2-Vision, on the CK12-QA dataset, evaluating their performance on both diagram-based and text-only questions. This includes a direct comparison of model performance across different input configurations, including standalone questions and retrieval-augmented settings, to assess the impact of contextual information. Furthermore, we conduct a quantitative analysis of the models’ ability to align multimodal questions with multimodal context. This work also proposes a retrieval-augmented evaluation framework for textbook question answering, which enables a systematic investigation of context-aware prompting strategies for MLLM-based answer generation. In addition, we identify catastrophic context interference, where adding retrieved context significantly degrades model performance on visual reasoning tasks. This highlights that more context is not always better, underscoring the need for improved context integration in MLLMs. Ultimately, this work advances our understanding of how MLLMs can be leveraged for complex educational reasoning tasks, highlighting both their potential and their current limitations in supporting curriculum-based learning.
@@ III Method
Given a query embedding $f(q)\in\mathbb{R}^{d}$, derived from the question text or image depending on the modality, and a set of candidate embeddings $g(c_{i})\in\mathbb{R}^{d}$ representing paragraphs or diagrams, we compute the cosine similarity to rank candidates, as shown in equation (1).
@@
All embeddings are generated using ImageBind [18], a cross-modal embedding model that projects inputs from various modalities, such as text and images into a shared embedding space. To store and index paragraph embeddings, we utilize KDB.AI [19], a high-performance, scalable vector database optimized for similarity search. During inference, the question embedding is used to query KDB.AI, which efficiently retrieves the top-$k$ context based on their cosine similarity to the query.
@@
Each input is composed of several elements to create a comprehensive prompt for the model. It includes a multiple-choice science question, an optional associated diagram such as a labeled process or structure, and a multimodal context consisting of a relevant paragraph and diagrams retrieved from the full set of CK12 lessons using a RAG approach.
@@
The proposed methodology preserves the original architecture of underlying pretrained multimodal large language models (MLLMs), requiring no structural modifications. Instead, it systematically integrates and further fine-tunes these existing MLLMs within a structured TQA framework. This approach enables task-specific adaptation while maintaining broad applicability and compatibility across diverse model architectures, thereby leveraging both the foundational strengths of pretrained models and the performance gains afforded by targeted fine-tuning.
@@
0p-0I, 1p-1I, 3p-3I = number of retrieved paragraphs and images; T-T/F = Textual True/False; T-MC = Textual Multiple Choice; All = overall accuracy across DQ and NDQ.
Experimental Setup
@@
Dataset
We utilize the CK12-QA dataset, a curriculum-aligned benchmark comprising comprising over 26,000 questions derived from the CK-12 Foundation’s open educational resources. The dataset includes textual true/false (T-T/F), textual multiple choice (T-MC), and diagram-based (Diagram Qs) items drawn from 1,076 lessons across Life Science, Earth Science, and Physical Science.
Input Conditions
Each model is evaluated under a zero-shot setting using two distinct input configurations to test its reasoning capabilities under different scenarios. The first is a ”Basic” condition, where only the question and its associated diagram, if applicable, are provided to the model. The second is a ”Multimodal RAG” condition, where a retrieved lesson paragraph is included alongside the question and its diagram to provide additional context.
Evaluation Metrics
@@
where $\hat{y}^{(n)}$ in equation (2) is the model prediction, and $\mathbb{1}[\cdot]$ is the indicator.
IV Results
IV-A Zero-Shot Performance
Tables III summarize validation and test accuracy for both models across varying input configurations. LLaMA demonstrates stronger baseline performance without additional context (e.g., 74.50% vs. 39.04% on T-MC in the 0p-0I setting–test set). However, LLaVA shows greater relative improvement when context is introduced. In the 3p-3I setting–test set, LLaVA’s T-MC accuracy improves by +14.16% (from 39.04% to 53.20%), narrowing the gap with LLaMA, whose performance declines by –17.11% (from 74.50% to 57.39%) on the test set. This indicates that LLaVA more effectively leverages retrieved textual context, whereas LLaMA’s performance is more sensitive to the presence of distractor paragraphs.
LLaMA significantly outperforms LLaVA on baseline diagram-based questions, achieving 74.07% accuracy on the validation set in the 0p-0I setting compared to LLaVA’s 46.89%. However, with added context, LLaMA’s performance on diagram questions–validation set degrades sharply, dropping to 25.93% in the 3p-3I setting, while LLaVA remains more stable at 42.89%. This suggests that LLaVA is more robust to multimodal interference introduced by retrieval-augmented input.
@@ IV-A Zero-Shot Performance
IV-B Fine-Tuned Model Results
During fine-tuning, we minimize the autoregressive cross-entropy loss as in equation (3):
@@ IV-B Fine-Tuned Model Results
TABLE III: Performance of fine-tuned LLaMA 3.2-Vision on CK12-QA under a 3 Paragraph, 3 Image configuration.
Metric Val (3p-3I) Test (3p-3I)
@@
TABLE IV: Performance of fine-tuned LLaVA on CK12-QA under a 3 Paragraph, 3 Image configuration.
Metric Val (3p-3I) Test (3p-3I)
@@
LLaVA shows a marked decline in overall performance following fine-tuning, with accuracy dropping from 48.92% to 27.78% on the validation set. Notably, the model exhibits inconsistent behavior across tasks in both test and validation sets, for example, while performance on T-T/F questions had a marginal drop on the test set (from 53.71% to 50.98%), accuracy on T-MC questions falls sharply to 19.24% from 53.20%, indicating instability in task generalization. Accuracy on diagram-based questions decreased across both validation and test sets, suggesting limited improvement in visual–textual alignment despite fine-tuning.
Fine-tuning outcomes diverge sharply between the two models. As shown in Table III, LLaMA 3.2-Vision demonstrates substantial improvements over its zero-shot performance, achieving 71.16% accuracy on the test set, up from 35.31%. Significant gains are observed across all task types: for example, on the validation set, T-MC accuracy increases from 59.15% to 79.76%, while performance on diagram-based questions rises from 25.93% to 65.88%, indicating effective multimodal learning. These results reflect strong cross-modal integration and suggest that fine-tuning successfully mitigates context interference present in zero-shot settings.
In contrast, LLaVA struggles to adapt effectively, as shown in Table IV. Its validation accuracy drops by nearly 50%, while LLaMA 3.2-Vision achieves 73.72% overall accuracy on the same split. This disparity highlights potential architectural advantages in LLaMA for multimodal fusion. LLaVA also shows task-specific instability: although T-T/F performance a slight decrease on the test and validation sets, T-MC accuracy declines markedly, pointing to inconsistent task prioritization. By comparison, LLaMA maintains balanced improvements across all tasks, suggesting more robust and curriculum-aware reasoning capabilities. Finally, whereas LLaMA effectively leverages retrieved context and resolves conflicts between textual and diagrammatic inputs, LLaVA appears to either overfit to noisy context or lack sufficient visual grounding, evidenced by its persistently low performance on diagram-based questions.
@@
A comparison of the models in the zero-shot setting reveals distinct behaviors when handling retrieved context. LLaVA shows a consistent ability to leverage this context, with its overall validation accuracy increasing from 44.45% in the basic setting to 48.92% when context is added, highlighting its capacity to integrate additional textual information for enhanced reasoning. In contrast, LLaMA 3.2-Vision, which achieves a much stronger baseline performance on diagram-based questions with 74.07% accuracy, experiences a sharp performance drop to 25.93% in the same setting when contextual paragraphs are introduced. This indicates a significant difficulty in fusing retrieved text with its visual reasoning.
IV-C Catastrophic Context Interference: A Deeper Analysis
The analysis of these results reveals two primary findings that are central. First, we identify a phenomenon we term ”catastrophic context interference,” particularly evident in the zero-shot performance of the more capable Llama 3.2 Vision model. While this model demonstrated strong baseline visual reasoning on diagram-based questions (DQs) without any external context (achieving 74.37% test accuracy), its performance plummeted dramatically to just 15.28% when a standard MM-RAG pipeline provided three relevant paragraphs and three images as context. This stark degradation highlights a critical failure mode: the naive application of a standard RAG pipeline, intended to provide helpful context, can be actively harmful to the model’s reasoning process for certain tasks. The model appears unable to reconcile or filter the retrieved textual information, allowing it to interfere with and override its otherwise strong visual grounding. This finding serves as a crucial cautionary insight for practitioners, demonstrating that more context is not always better and that the interaction between MLLMs and retrieved information is far from a solved problem.
It is important to note that our standard RAG pipeline, while robust, may be providing context that is topically relevant but not optimally salient for answering the specific question. This raises a key question: is the failure in the MLLM’s ability to ignore distracting information, or in the retriever’s ability to provide only essential information? A crucial direction for future work is to test a ”smarter” RAG that uses a re-ranking mechanism to more aggressively filter the context. This would help isolate the variable and more definitively test the MLLM’s context integration capabilities.
Second, the study reveals a stark divergence in fine-tuning adaptability between the two models. Llama 3.2 Vision, despite its zero-shot struggles with RAG context, showed a remarkable ability to learn from it during fine-tuning. Its overall test accuracy with RAG context improved from 35.31% (zero-shot) to 71.16% after fine-tuning, with performance on DQs recovering substantially. This suggests that its architecture possesses the latent capacity to learn the complex task of integrating and weighing information from multiple multimodal sources when provided with task-specific training data. In contrast, LLaVA 1.5, which showed more modest but stable zero-shot gains from RAG on text-only questions, exhibited a significant performance decline after fine-tuning. This suggests that its shallower modality fusion architecture may be more prone to overfitting on textual cues in the training data, leading to a loss of generalization and persistently poor performance on diagram-centric tasks. This goes beyond a simple ”model A is better than model B” conclusion, pointing instead to fundamental differences in architectural depth and capacity for multimodal learning.
IV-D Qualitative Error Analysis
To deepen our understanding of why catastrophic context interference occurs, a qualitative analysis of Llama 3.2 Vision’s errors in the zero-shot MM-RAG setting was conducted. This analysis revealed several patterns. In many cases, the retrieved text paragraphs, while semantically related to the question’s topic, were not directly relevant to answering the specific query posed about the diagram. For example, a question might ask to identify a specific part of the cell cycle in a diagram, while the retrieved text discusses the overall purpose of cell division in broad terms. The model would often seize upon keywords in the distracting text, leading it to an incorrect answer, even when the visual evidence in the diagram was clear.
Second, a common failure mode occurred when the retrieved text described a general scientific principle, but the diagram illustrated a specific instance or exception. The model frequently prioritized the general textual statement over the specific visual evidence, failing to correctly apply the general rule to the specific case shown in the diagram.
Additionally, While the RAG pipeline generally retrieved factually correct information, there were instances where the retrieved text might contain subtle nuances or a slightly different emphasis that conflicted with the simplified representation in the diagram, leading to model confusion. These examples illustrate that the interference is often not due to factually incorrect retrieved information, but rather due to the MLLM’s inability in a zero-shot setting to robustly prioritize sources of information (e.g., trust the specific diagram over the general text) and to identify the most salient piece of information from a noisy, albeit topically relevant, context.
IV-E Note on Retrieval Quality
The analysis presented assumes that the underlying MM-RAG pipeline, using ImageBind and KDB.AI for retrieval, provides context that is semantically relevant. A brief inspection of retrieval results confirms that the pipeline consistently retrieves paragraphs and diagrams that are topically related to the query. For instance, questions about the Earth’s core retrieve documents about geology and the Earth’s layers. Therefore, the observed ”catastrophic context interference” is not merely an artifact of poor retrieval but a genuine phenomenon of MLLM reasoning when presented with multiple, sometimes conflicting or distracting, pieces of relevant information. The core issue lies in the integration and utility of the retrieved context, not its topical relevance alone.
IV-F Discussion
@@
This discrepancy underscores the importance of context relevance filtering. Despite their strong vision-language capabilities, both models exhibit difficultly handling indiscriminately retrieved paragraphs. To mitigate this, more effective context selection strategies are needed, strategies that account not only for textual similarity but also for visual, textual alignment. Incorporating cross-modal attention mechanisms may offer a promising solution, enabling models to dynamically evaluate and prioritize context passages based on their relevance to both the question and associated visual input.